Back to the NIKA2 Pool page

Daily Reports

October 2, (Tuesday)

Andrew, Isabella, Jack, Laurent, Stefano, Matthieu and Bilal arrived at the telescope.

After a quick tutorial and a rough start, we manage to check pointing and focus on 2251+158. While it's bad, it's improving a lot with the night coming, and the weather is very promising.

We continue with a night of calibration measures, with beammaps on Uranus, Neptune, skydips and maps on the secondary calibrators, before we start with science projects.

[Andrew, Isabella and Jack took the shift Tuesday 10pm to Wednesday 6am.]

We started the shift with beam maps on Uranus and Neptune, followed by the @otf_loop_8_5 on each, before moving to NGC7027, CRL2688 and MWC349 for the otf loops.

After this, tau conditions appeared to be somewhat stable at tau~0.3, so we began science observations with the SZ LP 199-16, obtaining the 4 repetitions requested in the readme on PSZ2-G111.6. Up to this point, there had been some spikes in the tau, which was steadily trending upwards past 0.35. Next, after remaking the pointing and focus, we changed to project 094-18, and tried to make some maps of XLSSC102. The weather was becoming unstable, and trending about of 0.4, and we cancelled the observing script after 3/6 scans when it became clear that the conditions were no longer suitable for observing.

October 3, (Wednesday)

[Morning shift: Matthieu and Stefano]

The first hours were lost due to bad weather. Unstable and high tau (>0.5).

It became better after the sunset. When we restarted (10am local), we lost ~30min because the acquisiation software was down. Then, it was fixed by the AoD. Later, we managed to perform a focus and a skydip and started to observe the 083-18 program (PI:Oteo) until the end of the shift: J090910, J083932, J090613, J091454.

[Afternoon shift: Laurent and Bilal]

Nothing could be done until sunset. After that, we went for a few calibrating targets to prepare for the next shift...

[Andrew, Isabella and Jack took the shift Wednesday 10pm to Thursday 6am.]

As the shift begins, the opacity is ~0.4, with a downward trend and appears to be becoming more stable. We begin with a skydip, followed by beammaps and @otf_loop_8_5 on Neptune, then Uranus. As with the previous night, this was followed by a pointing and OTF loop on each of NGC7027, CRL2688 and MWC349.

At midnight UTC (2am local) after pointing and focus on 2251+158, we begin the first science observations of the day for project 199-16 . Although we completed 4/4 repetitions requested last night, we go back for another 2 since tau was higher than requested yesterday. This time the conditions are very similar, with tau of ~0.35, but appears to be stable as we take the 2 repetitions on PSZ2-G111.6. After a pointing and focus on Uranus, we moved on to project 094-18 and spent 3 hours on XLSSC102. Weather conditions continued to steadily improve during this time, closing in on tau~0.3

October 4, (Thursday)

[Morning shift: Matthieu and Stefano]

We replace the nightshift team at 6am local. They were executing a deep integration on an the XLSSC102 SZ cluster from the 094-18 program. We completed the integration to reach the integration time requested by the PI for each repetitions. We then observed a source from the 083-18 program (PI:Oteo) program. Then, the weather got better (tau~0.3 and stable), so we shifted to the observation of 070-18 as asked by the AoD. We observed JINGLE10 (the other sources were still low). Then, there was the sunrise (~8h15 local). So, we started a series of pointing and focus. We lost time because the acquisition crashed at 8h30 local (similar to the previous day). The AoD restart it. We made pointing,focus, and skydip, and then we started to integrate ~45min on GOODS-N (192-16). We performed a new focus to check the stability and got a bad 2mm beam (~19 arcsec). We had difficulty to obtain a good focus and the KIDS needed to be retuned. Ultimately, we managed to get a good focus and started to observe some JINGLE objects of the 070-18 program, since the weather was still good (AoD told us to stop integrating on deep field): JINGLE186, JINGLE122.

[Afternoon shift: Laurent and Bilal] The opacity is terrible at the beginning of the shift. Let's hope for the best for when the sun sets!

[Night shift: Andrew, Isabella & Jack]

We take over at 10pm local, with a beam map on Neptune underway, followed by a loop of 8x5 OTF scans. After a pointing on K3-50a, we begin science observations for project 050-18 (PI: Lestrade) on Vega with tau ~0.38 and improving rapidly. We spent from 11pm to 1am on this project, acquiring 3 repeats of the observing script, with tau finishing ~0.3.

At 1am, we returned to a session of OTF loops on NGC7027, CRL2688 and MWC349 for nikas-18. Next (2:15am), we moved onto project 094-18, for observations on the cluster XLSSC072 with a brief otf loop on Uranus following a focus (otf loop under nikas-18). In total, we made 2 repeats of the observing script on XLSSC072 over 2 hours.

At 5:18am, we moved on to project 091-18 to observe MS0735.6+7421. Tau has remained stable at 0.3 for most of this shift.

October 5, (Friday)

[Morning shift: Matthieu and Stefano]

After we took over after the night shift, we continued to observe 091-18. We then made a pointing+focus and observed GOODS-N (192-16) during one hour. The execution finished just before the sun rose above the mountain. The acquisition crashed just before the end of the last scan, which is only covered. This is the thrid time this week that it crashed at this exact time of the day. After Bilal restarted the acquisition, we made a pointing, focus, and then a skydip. We had much less difficulty than yesterday to get a good focus during the sunrise. So, we continued to observe GOODS-N for 2 extra hours. After that, we switched to 070-18. The weather started to degrade during the second integration of the day on JINGLE122, i.e. the 5/5 (tau went up to 0.5 with several spikes). Before this, the conditions were good for the season with tau~0.3 stable.

[Afternoon shift: Laurent and Bilal] As despair in the form of an everlasting rising opacity grows, we finish our shift with an improving weather observing for the duration of two scans the project 005-18.

[Night shift: Andrew, Isabella & Jack]

We take over the shift, tau is ~0.4 but rapidly improving. We begin with observations on Abell_2443 for project 090-18. At 00:20, we moved onto the calibrator sources MWC349, CRL2688, NGC7027, Neptune and Uranus for a series of @otf_loop_8_5. We finished the calibrators with a beam map on Uranus in good conditions (tau ~0.25). We moved onto 094-18 at 2:30am, and integrated on XLSSC072 for 1 hour. We then moved on to 095-18, and observed MOOJ0105 for approximately 1 hour. Tau had worsened in the previous set of observations, so we changed to project 094-19, afterwards and observed 091-18 for 50 minutes.

October 6, (Saturday)

[Morning shift: Matthieu and Stefano]

At the begining of the morning shift, we performed a focus and observed the 095-18 projet. We then observed ~1h45 GOODS-N (192-16). After the sun rose above the mountain, we performed a focus and a skydip. We then moved to the COSMOS field (192-16) for 2h. Finally, we observed 3 JINGLE sources (070-18) before the end of the shift. The weather was good and stable with tau~0.25. We did not have any particular technical problem (the acquisition did not crash around 9h contrary to the 3 previous days).

[Afternoon shift: Laurent and Bilal]

For once, the weather stays stable at the start of the afternoon. We observe project 005-18.

[Night shift: Andrew, Isabella & Jack]

The weather is lousy at the start of the shift. We start observing at 00:20 local time, when tau has dropped to ~0.4. At 01:00, tau is still ~0.4 but unstable, and we move to MWC349, CRL2688, NGC7027, Neptune and Uranus for a series of @otf_loop_8_5. After this, the weather became bad for a period, and we started observing for 091-18 (target MS0735) at ~4am. The tau is ~0.45, but Jack (co-I on this project) thinks this is worth a shot. We stopped at 5am due to conditions worsening again.

October 7, (Sunday)

[Morning shift: Matthieu and Stefano]

After a first hour without observation because of bad weather, we tried to observe when the tau started to decrease down to tau~0.5. We observed 091-18 during ~30 min, but the tau degraded again. We did a pointing and skydip at sunrise and stopped observing. The tau then continued to degrade and climbed up to ~0.65 at the end of the shift.

[Afternoon shift: Laurent and Bilal] Again, very bad conditions, so we stop observations. We will check from times to times if the weather is getting better. During the afternoon, as we were seeing the KIDs drift a lot, after retuning, we restarted the acquisition. Which would not restart. Cold reboot of crate, deactivation of individual boxes did not work. With Juan on Skype, Bilal had to go to the cabin and unplug crate 2, wait for a minute and plug it back again to have it back up. This situation was frightening, but ended up being solved quickly, thanks to Juan, Olivier and Alessandro's help!

[Night shift: Andrew, Isabella & Jack]

We take over the shift at 10pm, with very little observing having been done, with tau >> 0.4, and highly unstable. Conditions are now tau ~ 0.35, so we are starting observing with project 090-18 (target: Abell 2443). At 00:45 we start observing the calibrators MWC349, CRL2688, NGC7027 and Uranus. At 3am, we switched to project 091-18 for observations of MS0735.

October 8, (Monday)

[Morning shift: Matthieu and Stefano]

we take over the shift and continue the observations started at the end of the night shift (20 minutes on 091-18, source MS0735). After checking pointing and focus, we integrate further 80 minutes on the same source. At the end of the last repetition, the sun rises (above the horizon, but is still behing the mountain. Follow a skydip and the usual pointing/focus. The tau-meter marks now tau<0.3 and we observe 50+50 minutes the COSMOS field for the project 192-16. Finally, we switch to project 070-18 and observe source sJINGLE16 (one, missing repetiion) and JINGLE4 (four repetitions). During rep. 3 of JINGLE4, the tau-meter shows instabilities and spikes. A quick check on a nearby pointing source shows no sign of beam degradation, therefore we integrate further and complete the 4 repetitions requested by the P.I.

[Afternoon shift: Laurent and Bilal]

At the start of the shift, opacity increased to a very high value. We stop observation, and we will see during day how things goes.

[Night shift: Andrew, Isabella & Jack]

As we begin, conditions have just made a miraculous improvement to tau~0.25 after being awful all afternoon. We take a good chunk of data on Vega for program 050-18 (PI: Lestrade) until ~ 1:30am local time. At this point, the weather has become unstable, and so after finishing off the observing script, we change project to nikas-18 to make some OTF maps of the usual calibrators: NGC7027, CRL2688, MWC349 and Uranus. At 3am, with tau >~0.35, we move to project 091-18, to observe MS0735. At 4:15am, we started having trouble with focusing, and the conditions started looking bad. We were unable to focus on Uranus by 5 am, with tau oscillating wildly and so observations stopped there.


Overall, it was a very productive pool, compared to previous one (the Pool Manager and Matthieu still remember the March Catastrophe...). Thanks to all participants!

DailyReportsNika2Pool020918 (last edited 2018-10-09 10:31:14 by ladjelate)